### v2

# Dynamics of Anyon Clusters in Fractional Quantum Hall Fluids

Abstract
In fractional quantum Hall fluids, the quasiparticle excitations are anyons with fractional charges and statistics. Effective interactions among the anyons can be induced by either model or realistic electron-electron (e-e) interactions. Without losing the generality, we investigate such phenomena for the Laughlin $1/3$ and Moore-Read (MR) non-Abelian phases. Anyons display rich internal dynamics in both cases that will lead to interesting experimental consequences. In particular, bound states of two Laughlin anyons are preferred under short-range e-e interactions, leading to $2e/3$ – instead of $e/3$ – effective charge carriers at low temperatures, which have been seen in several experiments. MR phases host two topologically distinct fusion channels: “1” and $\psi$”. The different effective interactions of $e/4$ anyons in the two sectors suggest the vanishing of the degeneracy of fusion channels when the e-e interaction is no longer its model Hamiltonian, in which case different bound states could also appear. This indicates the possibility of energetically manipulating the two types of anyons by tuning the bare e-e interactions. We point out that the results of recently developed high-resolution STM measurements will be affected by the effective anyon interactions, where anyons are clustered together after the tunneling of electrons. The low-lying parts of the local density of states affected by various anyon clusters are simulated for both Abelian and non-Abelian systems with (screened) Coulomb interactions.
I Introduction
Anyons with fractional charges and fractional statistics are particles only existing in two-dimensional manifolds because of the nontrivial braid group in 2D [1, 2, 3, 4, 5, 6]. The fractional quantum Hall (FQH) states, realized in a two-dimensional electron gas (2DEG) with strong electronic interaction subjected to a perpendicular strong magnetic field at low temperature, provide an ideal platform to study the exotic properties of anyons [7, 8, 9, 10]. Quasiholes that can be regarded as anyons in the incompressible FQH fluids are low-lying collective excitations of electrons created by adiabatically inserting fluxes into the FQH ground states with distinct topological orders [11, 12]. The simplest example is the Laughlin $1/m$ anyons with fractional statistics $\nu\pi$ and fractional charge $\nu e$, created by inserting one flux into the Laughlin ground state at filling factor $\nu=1/m$ [13]. More intriguing cases are the non-Abelian phases in which several anyons with non-Abelian statistics are created by one flux insertion. The great interest in non-Abelian anyons is rooted in the possibility of their application to fault-tolerant topological quantum computing [14, 15, 16].
Much effort has been made to understand the properties of emergent quasi-particles in FQH fluids, such as their fractional statistics, the relationship between them in different FQH phases, the size of anyons and their structures, etc [17, 18, 19, 20, 21, 22]. One modern perspective is that anyons are “elementary particles” within the respective conformal Hilbert spaces (CHSs) defined by model Hamiltonians, or more generally, by local exclusion conditions (LEC) [4, 23, 24, 25, 26, 27]. These CHSs with emergent conformal symmetry are spanned by the ground state and all quasihole states [28, 29]. If only the model Hamiltonian defining the CHS exists as the electronic interaction term, then the anyons are massless (i.e., it costs no energy to create a single anyon) and free (i.e., the anyons are non-interacting) particles. With more realistic electron interactions, they will gain self-energy and start to interact [30]. Previous studies manifest the hierarchical structure of CHSs, implying the nontrivial internal structures of anyons as the elementary particles in the corresponding CHSs [28, 31]. These intricate properties and the nature of CHSs serve as the foundation for revealing the exotic properties of different topological FQH phases and their relationship [28, 20, 29, 32, 33].
Experimental progress, such as shot noise, quantum point contact, and anyon interferometry, provides direct detection for both Abelian and non-Abelian natures of certain anyons [34, 35, 36, 37, 38, 39, 40]. While these experimental methods often require a dilute anyon gas for braiding and quantum tunneling, meaning a large separation between anyons, other phases such as Wigner crystal, stripes, and bubbles will compete with FQH phases when the anyon’s density is high [41, 42, 43, 44, 45]. This leads to the rich phase diagrams of 2DEG and makes it interesting to understand the dynamics of a cluster of anyons when they are close and their interaction becomes important. Recent advancements in atomic-scale spectroscopy within ultra-clean, diverse 2D systems have enabled the use of local scanning tunneling microscopy (STM) to measure tunneling processes of a cluster of anyons, probe anyon clusters’ dynamics, and provide valuable insights into their energies and fusion rules [46, 47, 48, 49]. In such experiments, the proximity of anyons amplifies the role of effective interactions that strongly influence their behavior [50].
In this work, we focus on the effective microscopic interaction between anyons in the Laughlin and Moore-Read (MR) phases, derived from both the model and the realistic bare interactions between electrons. Such effective interactions are highly nontrivial, reflecting information about the particle statistics[3, 7] and the spatial extension of particle sizes (i.e., anyons are not point particles)[51, 22]. This leads to the formation of few-anyon clusters and are potentially important for many-anyon quantum fluids with both universal and non-universal dynamics. In particular, we show that two Laughlin 1/3 anyons form a bound “molecule” with short-range electron-electron (e-e) interactions, providing a possible explanation for the anomalous results in quantum point contact experiments of Laughlin phases [52, 53, 54, 55]. Moreover, different dynamics of non-Abelian $\psi$-type and Abelian $1$-type anyons in the Moore-Read phase with the same e-e interactions break the ground state degeneracy, implying that the Majorana fermion is no longer massless in the presence of the realistic interactions. Here the term ”dynamics” is used in the general sense for the parametric evolution of quasihole clusters’ energies, which will manifest themselves in various experiments where dynamical properties of the quantum Hall states are measured. For both Abelian and non-Abelian anyons, we demonstrate that the local density of states (LDOS) detected by high-resolution STM depends on the interactions within the anyon clusters created by adding or removing a single electron. This makes STM a powerful tool for probing the effective anyon interactions in topological FQH phases.
This paper is structured as follows. Section II first briefly reviews the formalism of the microscopic theory for FQHE on spherical geometry, including Haldane pseudopotentials and the Jack polynomial method. Then the dynamics of two- and three-anyon clusters of the Laughlin 1/3 kind induced by both model and realistic electron-electron interactions is studied. Similarly, in section III, the dynamics of two- and four-anyon clusters for the MR kind is investigated, and the distinct interacting behaviors of the two types of MR anyons in systems with odd and even electron numbers under the same electron-electron interaction are highlighted. Finally, in IV, we explore the potential experimental signatures of the anyon dynamics, particularly in high-resolution tunneling measurements.
II Interactions within a cluster of Laughlin quasiholes
II.1 Spherical Geometry and Jack Polynomial Formalism
[…]
Assuming the electrons are fully spin-polarized, an isotropic two-body interaction in the $n_{th}$ LL, neglecting LL mixing, can be written as
[…]
$\psi_{(\alpha,\beta)}^{l,m}=(u_{1}v_{2}-u_{2}v_{1})^{m}\prod_{i=1,2}\left(\bar{\alpha}u_{i}+\bar{\beta}v_{i}\right)^{2l-m}$ (2)
[…]
The projected interaction can then be decomposed as:
$\hat{V}^{(n)}=\sum_{m=0}^{2l}c^{(n)}_{m}\hat{V}_{m}$ (3)
[…]
$|\hat{V}_{m}=\sum_{M=-L}^{L}\ket{L,M}\bra{L,M}$ (4)
is the PP operator, which projects any two-particle state onto the subspace of relative angular momentum $m$. Unless otherwise stated, all analyses in this paper are performed in the lowest LL (LLL), corresponding to the strong magnetic field limit, and the LL index $n$ is omitted hereafter. PPs provide both an intuitive physical picture and a sparse representation of the interaction, thereby simplifying numerical calculations. More importantly, they define model Hamiltonians whose eigenstates are the model wavefunctions of certain FQH phases. For instance, the eigenstates of the model Hamiltonian $\sum_{k}^{p}\hat{V}_{2k-1}$, $k=1,2,$..., are the fermionic Laughlin model wavefunctions, describing both the ground states and anyonic quasihole states at the filling factor $\nu=1/(2p+1)$. These states constitute the null space of the model Hamiltonian $\sum_{k}^{p}\hat{V}_{2k-1}$ and span the corresponding conformal Hilbert space (CHS). Understanding the effects of individual PPs can also help us predict the influence of realistic interactions that are linear combinations of PPs.
Jack polynomial formalism expresses the many-body wavefunctions of FQH states by root configurations with different admissible rules[62, 63]. Root configurations consist of “1” or “0”, each representing a single-particle orbital on the sphere with azimuthal angular momentum s=l,l+1,,ls=-l,-l+1,...,l from the leftmost (north pole of the sphere) to the rightmost (south pole of the sphere). “1” indicates an orbital occupied by one electron, whereas “0” represents an unoccupied orbital. Laughlin 1/3 states satisfy (1,3) admissibility rule, meaning that there is no more than 1 electron in 3 consecutive orbitals. The ground state for Laughlin 1/3 conformal Hilbert space $\mathcal{H_{L}}$ is the highest density state[64]:
10010010010011001001...001001 (5)
where the number of orbitals $N_{o}=3N_{e}-2$, $N_{e}$ being the electron number. Electrons in the ground state symmetrically occupy the orbitals, $L=L_{z}=0$. All the other quasihole excitation states in $\mathcal{H_{L}}$ can be obtained by inserting fluxes ($0s$) into the ground state.
In Abelian phases such as the Laughlin states, one anyon is created by each flux insertion, namely the number of quasiholes $N_{qh}$ equals the number of extra fluxes. The new orbital indices after the insertion will be $s^{\prime}=-(l+N_{qh}/2),-(l+N_{qh}/2)+1,...,l+N_{qh}/2$. When all fluxes are inserted at the north pole, each electron’s spin $s$ will increase $N_{qh}/2$, yielding the state with total angular momentum $L=L_{\text{max}}=N_{qh}N_{e}/2$, such as Eq.(6) in which two more fluxes are added. Empty circles below the digits denote the positions of anyons. We define the relative angular momentum of an anyon cluster in a state with total electron angular momentum $L$ as: $\Delta M=L_{max}-L$. Pinning one flux at the north pole and moving another away increases $\Delta M$ as the anyons separate, shown by the examples in Eq.(7). If the two fluxes are added at the north and south poles, the electron occupation remains symmetric as in Eq.(8), resulting in $L_{z}=0$ and $\Delta M_{max}=L_{max}$
 0  0 $1\displaystyle\hbox{\hskip 4.16663pt\hskip-2.5pt\hbox{$0$}\hskip-2.5pt\hskip-4.16663pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\circ$}}\hskip-4.16663pt\hskip 4.16663pt}\hbox{\hskip 4.16663pt\hskip-2.5pt\hbox{$0$}\hskip-2.5pt\hskip-4.16663pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\circ$}}\hskip-4.16663pt\hskip 4.16663pt}100100$...1001 (6)
[…]
Figure 1: (a) Interaction between two anyons with model electron-electron interaction $\hat{V}_{3}$ (blue circles) and $\hat{V}_{5}$ (purple squares) in $\mathcal{H_{L}}$; (b) Effective interaction between two anyons induced by bare electron-electron interactions: Coulomb (olive squares) and Yukawa with $\lambda=0.375/l_{B}$ (green circles), inset: energy differences between the bunching state $E_{\Delta M=2}$ and the most separated state $E_{S}$ ($\Delta M=12$ for the largest system size $N_{e}=12$ we computed) against $\lambda$. (c, d) Finite-size scaling of the energy differences between the most separated state $E_{S}$ (orange crosses), or the bound state $E_{B}$ when $\Delta M=0$ (yellow pluses), and the bunching state $E_{\Delta M=2}$ with $\hat{V}_{3}$ (c) and $\hat{V}_{5}$ (d).
II.2 Dynamics of Two Laughlin 1/3 Anyons
We first study the interaction between two Laughlin 1/3 anyons under several leading two-body PPs using the model Hamiltonian
$\hat{H}_{\text{int}}=t\hat{V}_{1}+\hat{V}_{2k+1},\ k=1,2,3,$... (9)
with $t\gg 1$, projecting the studied Hilbert space to $\mathcal{H_{L}}$ (null space of $\hat{V}_{1}$) where the elementary particles are the Laughlin $1/3$ anyons, undressed from neutral excitations outside of $\mathcal{H_{L}}$. In our calculations, $t=100$ is typically set, which is sufficient to numerically enforce the projection, shown by a finite energy gap between the quasihole states and all the other states above them in the spectrum. In fact, we find that a value of $t$ just one order of magnitude larger than the coefficient of the additional e-e interaction term already ensures an effective separation between the quasihole sector and the higher-energy continuum.
Without $\hat{V}_{2k+1}$, these anyons are massless and non-interacting as an ideal gas in $\mathcal{H_{L}}$. After adding a relatively small $\hat{V}_{2k+1}$, anyons will gain mass and start to interact. Figure 1 (a) illustrates the effective interaction between two Laughlin $1/3$ anyons as a function of $\Delta M$ under the influence of the model electron-electron interactions $\hat{V}_{3}$ and $\hat{V}_{5}$. The interaction energy is defined as the total energy subtracted by the ground-state energy and the creation energies of all the quasiholes.
$E_{\text{int}}=E_{\text{total}}-E_{\text{g}}-N_{qh}E_{\text{qh}}$ (10)
$E_{g}$ stands for the ground state energy and $E_{qh}$ is the self-energy of one quasihole (energy cost to create one quasihole). To compute $E_{qh}$, we introduce one additional flux into the system while keeping the number of electrons fixed, and subtract the resulting ground-state energy from $E_{g}$. This defines the creation energy of a single, well-isolated quasihole, independent of any angular momentum, spatial configuration, or presence of other quasiholes.
The interacting energies between two anyons are negative at a relatively small separation under both $\hat{V}_{3}$ and $\hat{V}_{5}$ electron-electron interactions in $\mathcal{H_{L}}$, shown in Figure 1. Consequently, the two anyons will combine to form molecule-like structures with $\Delta M=2$ or $4$ being the lowest energy states, respectively. Finite-size scalings in (c, d) demonstrate that the key features, such as the formation and stability of the bound anyon cluster state, are robust and converge well when increasing system size. While a detailed understanding of the effective interaction’s origin remains to be established, our studies and several indirect arguments in the literature support the relevance to the non-trivial statistics and real-space density oscillations of the anyons [65, 22, 19]. A deeper investigation into the precise roles of these features is an open question for future work.
Figure 2: Energy spectrum of three-quasihole Laughlin 1/3 state with electron-electron (a) Coulomb interaction (b) Yukawa interaction when $\lambda=1/l_{B}$, both in the LLL; (c) Energy differences between the $\Delta M=6$ state and its competing ones-$\Delta M=8$ (olive circles) and $\Delta M=12$ states (green squares)-plotted as a function of $\lambda$, both exhibiting monotonic behaviours. Red crosses indicate quasihole states, blue pluses denote other excitations outside the QH manifold. The system size is $N_{e}=10$.
Building on the study with model Hamiltonians, we further explore the impact of realistic e-e interactions on two-anyon clusters. Forming an anyon molecule requires a relatively large coefficient for $\hat{V}_{1}$ and the dominance of $\hat{V}_{2k+1}$ with small $k$ based on the results of model Hamiltonians. We hence examine the anyon dynamics under the short-range Yukawa interaction that is commonly used to incorporate the screening effect theoretically:
$V(\mathbf{r})=\frac{e^{-\lambda|\mathbf{r}|}}{|\mathbf{r}|},\ V(\mathbf{q})=\frac{1}{\sqrt{\lambda^{2}+|\mathbf{q}|^{2}}}$ (11)
where $\lambda$ is the inverse of coherent length in units of $1/l_{B}$. Two Laughlin $1/3$ anyons tend to separate with Coulomb interaction, shown in Figure 1 (b). As $\lambda$ increases, the lowest energy state of the two-anyon cluster undergoes a monotonic transition from a deconfined state to a bunching state with $\Delta M=2$, such as the green dashed line manifested where the effective anyon interaction is induced by $\lambda=0.375$ Yukawa interaction. The inset of Figure 1 (b) shows the monotonic trend of this transition, where the energy differences between the bunching $\Delta M=2$ state and the separated state decrease gradually as $\lambda$ increases, changing from positive to negative. Since screening uniformly reduces the overall energy scale, we normalize the pseudopotential decomposition by fixing the coefficient of $\hat{V}_{3}$ to unity, and assign the remaining pseudopotential coefficients according to their original relative ratios. This ensures that all quantities such as energy gaps and total energies remain comparable across different values of $\lambda$ (and Coulomb).
The two-anyon “molecule” introduced by the short-range electronic interactions provides a plausible explanation for the unexpected results in quantum point contact experiments, where the effective charge carrier shows a temperature dependence, transitioning from $2e/3$ at low temperatures ($\sim 10mK$) to $e/3$ at higher temperatures ($\sim 100mK$) [52, 53, 54, 55]. In shot noise experiments, Coulomb screening will arise from the 2D dielectric environment, nearby metallic gates, and quantum point contact potentials, which will distort electric fields, suppress long-range interactions, and enhance short-range effects [66, 67, 68]. Based on the above discussion, at near absolute zero temperature with the effect of short-range interaction, the effective charge carriers in the quantum fluid are bunching “molecules” with $e^{*}=2e/3$ formed by two Laughlin 1/3 anyons. Thermal fluctuations will gradually break apart these “molecules” as temperature rises, and the charge carrier will completely change back to the single Laughlin 1/3 anyons with $e^{*}=e/3$ in the end. A quantitative comparison between the binding energy of the anyon “molecule” under Yukawa interaction with $\lambda=0.375/l_{B}$ ($10^{-24}\sim 10^{-25}\text{J}$) and the thermal energy at $T\sim 10-100\text{mK}$ confirms that their scales are comparable, supporting this scenario.
II.3 Interaction Within Three Laughlin 1/3 Anyon Cluster
To prepare for our later discussions about bulk tunneling experiments—which is essential for understanding many-anyon FQH fluids in realistic scenarios—we investigate the interactions within a cluster of three Laughlin 1/3 anyons. This configuration naturally arises when one electron is removed from the Laughlin 1/3 state. Although the three-anyon case is more complex than the two-anyon scenario, it follows a similar trend. The total relative angular momentum reflects the extent of the anyon cluster. The ground states of both $\hat{V}_{3}$ and $\hat{V}_{5}$ in $\mathcal{H_{L}}$ correspond to a three-anyon cluster with total relative angular momentum $\Delta M=6$. The interacting energy as a function of the total relative angular momentum $\Delta M$ within this anyon cluster is provided in the Supplementary Materials, where finite-size scaling of the interaction energy gaps between $\Delta M=6$ and its competing states in both cases confirms the stability of $\Delta M=6$ state in the thermodynamic limit.
We studied the three-anyon spectrum for both Coulomb and Yukawa potentials with varying screening length $\lambda$. In all panels of Figure 2, red crosses mark states in the quasihole CHS, identified via a large-t projection using the model Hamiltonian $\hat{H}=t\hat{V}_{1}+\hat{V}_{\text{int}}$, where there exists a finite energy gap between the quasihole manifold and the other higher-energy states. They are tracked as $t\rightarrow 0$ toward the physical interaction, during which the structure of the QH manifold – particularly its state counting – remains unchanged because of its topological nature [69]. Blue plus symbols denote all the other states outside the quasihole manifold, which are not the subject of this paper and won’t be further identified.
With LLL Coulomb interaction, the three-anyon cluster’s lowest energy state is $\Delta M=12$, indicating a loosely bound structure. However, the configuration transitions to a more compact cluster with $\Delta M=6$ as $\lambda$ in the Yukawa potential increases, corresponding to shorter-range electron-electron interactions. This transition is shown by depicting the energy difference between $\Delta M=6$ ($E_{\Delta M=6}$) and its competing states $\Delta M=8$ or $\Delta M=12$ as a function of $\lambda$ in Figure 2 (c). $\Delta M=6$ gradually becomes the lowest energy state as $\lambda$ increases. This aligns with our earlier finding in two-anyon cases that sufficiently short-range interactions dominated by $\hat{V}_{1}$ can replicate the effects of the model Hamiltonian Eq.(9) on the effective anyon interaction. It can also be seen from the whole spectrum of the system that the quasihole excitations (red crosses) gradually separate from other states (blue pluses) when increasing the $\lambda$, since during this process, $\hat{V}_{1}$ gradually becomes more dominant. We estimate the size of the anyon cluster using the first and second moments of their density distribution [21]. The estimated spatial radius of the lowest-energy anyon cluster is $40\sim 60\text{nm}$. Density for the system under Coulomb e-e interaction indicates long-range correlations and oscillatory structures, while for Yukawa interaction, it has a prominent central peak and mild, smooth oscillations extending toward the edges, reflecting strong short-range binding  [70].
III Interactions within a cluster of Moore-Read Anyons
The Moore-Read (MR) FQH fluid has attracted much greater attention because of its non-Abelian nature and potential applications in quantum computations. Unlike the case of the Laughlin state, two quasiholes, each with a $e/4$ charge, will be created by one single flux insertion in the MR ground state because of the flux fractionalization in non-Abelian phases. When fused, the two-quasihole excitation with $e/2$ charge collectively carries topological quantum numbers “11” or $\psi$”, exhibiting a statistical phase 0 or $\pi$ respectively [17]. The topological numbers of these two kinds of excitations come from the fact that they follow the Ising model in conformal field theory (CFT) with energetically degenerate fusion channels “11” and $\psi$”, corresponding to the U(1) boson and Majorana fermion chiral edge modes [4, 71, 15, 72]. For the same reason, a single $e/4$ quasihole excitation corresponds to the index $\sigma$” in the Ising CFT model. In a general MR phase, the “11” and $\psi$” types of quasihole excitations coexist, which is the reason for non-Abelian statistics. The degeneracy of a MR phase containing $N_{qh}$ quasiholes is $2^{N_{qh}/2-1}$.
III.1 Dynamics of Two MR Anyons
The root configuration for the MR ground state is 1100110011110011...0011, satisfying the $(4,2)$ admissible rule and the condition $N_{o}=2N_{e}-2$. We begin with the case of two quasiholes (one extra flux quanta) where the topological number of the system can only be either 1 or $\psi$, depending on whether the number of electrons ($N_{e}$) in the finite system is even or odd. These two types of two-quasihole excitations are illustrated by the following root configurations:
 1 $\hbox{\hskip 6.66664pt\hskip-2.5pt\hbox{$\color[rgb]{1,1,1}{1}$}\hskip-2.5pt\hskip-6.66664pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\circ\circ$}}\hskip-6.66664pt\hskip 6.66664pt}01100110011...110011$ 11-type, even $N_{e}$ (12)
 1 $\hbox{\hskip 6.66664pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-6.66664pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\circ\circ$}}\hskip-6.66664pt\hskip 6.66664pt}00110011...11001$1 $\psi$-type, odd $N_{e}$ (13)
Each empty circle represents one quasihole with $e/4$ charge. The parity of the number of electrons in between the pair of $e/4$ MR anyons is the same as the parity of the whole system’s $N_{e}$.
To compute the self-energy of a single $e/4$ quasihole in the MR phase, we fix the two quasiholes at the north and south poles of the sphere by Jack polynomials as 1010101011010...10101 and extract the energy from the ground-state energy $E_{g}$ with the same electron number. In the large-system limit, where their mutual interaction vanishes, the residual energy should approach twice the quasihole self-energy. More detailed explanations and discussion are provided in the Supplementary Material.
Figure 3: Effective interaction of two MR quasiholes with different electron-electron interactions. (a) $\hat{V}_{1}$ and (b) $\hat{V}_{3}$ within $N_{e}=19$ (odd, pink stars) and $N_{e}=20$ (even, blue squares) systems, calculated from ED; (c) Coulomb interaction and (d) Yukawa interaction when $\lambda=0.5/l_{B}$ in systems of $N_{e}=51$ (odd, pink) and $N_{e}=50$ (even, blue), computed by Monte Carlo with error bars shown.
First, the effective interaction between two MR anyons under the model e-e interaction is studied. Haldane pseudopotentials $\hat{V}_{1}$ and $\hat{V}_{3}$ are diagonalized within the null space of the MR model Hamiltonian, defined by a special three-body interaction[73, 24] and obtained by a similar strategy as Laughlin’s case Eq.(9) numerically, with the presence of two quasiholes. The effective anyon interactions defined by Eq.(10) with $N_{qh}=2$ are shown in Figure 3 (a) and (b), plotted against their relative angular momentum $\Delta M$. It depends on the parity of the system, as a manifestation of the odd-even effect in the MR phase [17]. Since the resulting dynamics have distinct behaviors under different PPs by which the e-e interaction can always be decomposed, it suggests that tailored electron-electron interactions could exploit this distinction to manipulate the two types of MR anyons. For instance, the e-e interaction dominated by $\hat{V}_{1}$ induces the effective interactions for $\psi$-type anyons to bind together while favoring a comparatively large structure of $1$-types. Conversely, to introduce an effectively repulsive interaction between the $\psi$-type anyons for non-abelian braiding (therefore the possibility of topological quantum computation), e-e interaction dominated by $\hat{V}_{3}$ is preferred.
Figure 4: Spectrum of MR with four quasiholes under various realistic interactions, with odd (here $N_{e}=13$, first row) and even (here $N_{e}=14$, second row) numbers of electrons. (a) Coulomb interaction; (b) Yukawa with $\lambda=0.5$; (c) Yukawa with $\lambda=1$.
Effective interactions between two MR anyons under realistic electron-electron interactions are more challenging to study due to the strong finite-size effect, which requires larger system sizes that are difficult to achieve by exact diagonalization. To overcome this limitation, we employ the Monte Carlo simulation to explore larger systems. The energies when two localized quasiholes are gradually separated in systems with $N_{e}=50$ (even) and $N_{e}=51$ (odd) are computed using the Pfaffian wavefunction by fixing one anyon at the north pole and moving the other apart. The final average stable value, corresponding to the regime where the anyons are sufficiently separated and no longer interact, is taken as the background energy and subtracted from all data points.
The resultant effective interactions between two localized $e/4$ MR quasiholes are plotted as a function of their chord distance. The odd-even effect remains prominent, with nearly opposite dynamical characteristics observed between systems with the two different parities. Under Coulomb interaction shown in Figure 3 (c), the effective interactions are repulsive at all distances for systems with $N_{e}$ of both parities. With shorter-range Yukawa interactions example $\lambda=0.5/l_{B}$ in Figure 3 (d)-the effective interaction for $\psi$-types is attractive at small separations. But it remains repulsive for $1$-types with the same distances, then develops to be attractive at a comparatively large separation. This behaviour indicates that short-range interactions energetically favor tightly bound pairs of $e/4$ MR anyons in odd $N_{e}$ systems, while promoting spatially extended, loose configurations of them in even $N_{e}$ systems. It means that the $\psi$ fusion channel is energetically more favorable under the short-range interaction than the $1$ channel when considering the fusion of two $\sigma$” anyons with $e/4$ charges. These results further support the feasibility of engineering bare electron-electron interactions to selectively separate and control the two types of fusion channels of the $e/4$ MR anyon and the consequent $e/2$ excitations.
III.2 Interaction Within Four MR Anyon Clusters
We now look at a cluster of four MR quasiholes—each carrying a $e/4$ charge—which will be useful in Sec . IV. A four-MR-anyon cluster emerges upon the removal of a single electron from the ground state. The resulting system exhibits a degeneracy of $2^{4/2}-1=2$, where $1$- and $\psi$-types could coexist, adding significant complexity to the analysis. Specifically, the two pairs of quasiholes can only be two 1-types or two $\psi$-types in even $N_{e}$ systems, whereas the configuration necessarily consists of one 1-type and one $\psi$-type in odd $N_{e}$ systems [74]. The energy spectrum of this system under model electron-electron interactions $\hat{V}_{1}$, $\hat{V}_{3}$, and $\hat{V}_{5}$ in the MR null space (in supplementary [70]) reveals two key features: a persistent odd-even effect in systems with four MR quasiholes and a preference of loose anyon configurations characterized by large relative angular momentum $\Delta M$ within the four-anyon cluster.
To extend our investigation to more realistic scenarios, given that MR states are proposed to be stabilized in the 1LL, we studied the e-e Coulomb and Yukawa potentials in the 1LL shown in Figure 4. In contrast to Laughlin’s case, the ground states of MR systems with four anyons consistently exhibit a loose structure and remain largely insensitive to variations in electron-electron interactions. For the system with $N_{e}=13$, the lowest energy state is the one with relative angular momentum $\Delta M=12$, the maximum allowed $\Delta M$ value in this finite system. In the $N_{e}=14$ case, the four-anyon cluster attains the largest possible value of $\Delta M=14$ under the 1LL electron-electron Coulomb interaction, and changes to $\Delta M=12$ for both Yukawa interactions we computed with $\lambda=0.5$ and $1$. The results are not conclusive due to the limited system size we can reach. Another observation is that the quasihole excitation spectrum maintains a comparatively robust, model-independent structure, unchanged not only across the parameter regimes studied among the realistic interactions, but also across the realistic interaction and the model Hamiltonian inside the MR CHS, whose spectrum can be further found in Supplementary Materials. Using the first and second moments of the density distribution, the estimated real-space radius of the four-quasihole cluster is $100\sim 140\text{nm}$ for the lowest energy state as $\Delta M=12$ in $N_{e}=15$ (odd) system. And for the $\Delta M=14$ anyon cluster in $N_{e}=16$ (even) system, it is $75\sim 110\text{nm}$. The spatial extent of the anyon clusters with the lowest energy in the MR phase exhibits a difference between odd and even electron systems, suggesting a parity effect in the cluster configuration in the finite systems. And the predicted spatial extents are much larger compared to Laughlin’s cases  [70].
IV Bulk tunneling experiments
The effective interactions within anyon clusters have direct experimental relevance in recently developed STM measurements, whose detailed setups and implementations, both theory and experiment, have been discussed in several recent works [46, 49, 48]. In such experiments, single-electron tunneling occurs when the electronic potential $V_{t}$ applied to the tip satisfies
$E=eV_{t}\geq E_{0}+E_{z}+E_{xy}$ (14)
where $E_{0}$ is the work function of the 2DEG, $E_{z}$ accounts for the finite-thickness effect [75], and $E_{xy}$ represents the energy scale of the strongly correlated 2D system studied here. After the bulk tunneling, a few-anyon cluster with total positive charge of +e+e will be left behind. The number of anyons in this cluster is determined by the specific FQH phase (e.g. $3$ for Laughlin 1/3 and $4$ for MR). Due to the non-trivial inter-anyonic interactions, as we have shown in the previous sections, the anyon cluster could be in different forms of “bound states” depending on the effective anyon interaction, the size of the STM tip, and the bias voltage applied to the tip. In particular, the effective interactions and the consequent structures of the few-anyon clusters influence the experimental outcomes, especially for the low-lying parts, which are mainly contributed by the quasihole states. It makes such measurements potentially very useful for understanding the dynamics of anyon clusters.
Figure 5: Possible low-energy LDOS for Laughlin 1/3 state with $N_{e}=10$ at FWHM = $0.003\text{(first panel)}/0.005\text{(second panel)}\ e^{2}/\epsilon l_{B}$. The entire background energies are subtracted in the subfigures since only the energy differences matter here. (a) Coulomb interaction, (b) Yukawa interaction when $\lambda=1$, both in the LLL.
The local density of states (LDOS) is a key observable in STM measurements. High-resolution tunneling can reveal a clear gap of approximately $0.1meV$, which corresponds to $\Delta E=0.002\sim 0.006\ e^{2}/\epsilon l_{B}$ whose exact value depends on the specific dielectric constant $\epsilon$ and the magnetic field $B$ under detection. According to this range, we simulated the possible LDOS for the 3-anyon Laughlin 1/3 states and 4-anyon MR states, focusing on the low-lying parts of the states mainly from the quasihole manifold. It is found that the variations in effective anyon interactions—induced by different bare e-e interactions—lead to distinct LDOS patterns, including differences in the numbers and positions of the peaks.
In this work, we follow the standard STM approximation where the tunneling amplitude is assumed to be constant and featureless, so the measured tunneling conductance is proportional to the local density of states (LDOS) of the sample. This approach omits the energy-dependent tunneling matrix element computed in more microscopic treatments such as Bardeen’s formalism, which derives from Fermi’s golden rule[76]. While the detailed tunneling amplitudes may vary with the STM tip geometry, material-specific parameters, or symmetry constraints, our analysis aims to capture the qualitative LDOS features arising from the internal structure of few-anyon bound states. A more refined modeling of the tunneling matrix elements will be pursued in future work.
Our analysis shows that the low-lying tunneling spectrum is rich, affected by the effective anyon interaction and the internal structure of the anyon cluster, both in the Laughlin and Moore-Read phases. We only present a preliminary study here, assuming the bias voltage of the STM tip does not alter the anyon clusters’ densities. This may not be the case given that the one-body potential for an electron to tunnel highly depends on the geometry of the tip, the bias voltage, as well as the distance between the tip and the Hall manifold [74]. A more detailed study with explicit experimental parameters will be presented in the future.
IV.1 LDOS Measurement for Laughlin 1/3 Phase
The low-energy spectrum for a system containing three Laughlin 1/3 anyons with realistic electron-electron interactions has already been shown in Figure 2 (a), where the most tightly bound anyon cluster state ($\Delta M=0$ state) lies at significantly higher energy than the other low-lying excitations for relatively long-range e-e interactions such as LLL Coulomb. As the screening parameter $\lambda$ increases, all quasihole excitations shift downward and gradually separate from the high-energy continuum. The $\Delta M=0$ state eventually becomes distinguishable from the gapped excitations but remains the highest in energy among all quasihole states. The simulated LDOS for $FWHM=0.003$ and $0.005$ under e-e Coulomb interaction and $\lambda=0.5$ Yukawa interaction, both in LLL, are shown in Figure 5. The LDOS results for different interactions are clearly distinguishable as in Figure 5. This indicates that LDOS can serve as a fingerprint for characterising the effective interactions among anyons.
Figure 6: LDOS of the MR system containing 4 anyons at FWHM = $0.001\ e^{2}/\epsilon l_{B}$ for: 1LL Coulomb interaction with (a1) $N_{e}=13$ and (a2) $N_{e}=14$; 1LL Yukawa interaction when $\lambda=0.5$ with (b1) $N_{e}=13$ and (b2) $N_{e}=14$.
IV.2 LDOS Measurement for Non-Abelian Moore-Read Phase
In the MR phases, the $\Delta M=0$ state is even more deeply buried in the continuum of gapped excitations, regardless of the e-e interaction being long-range 1LL Coulomb or comparatively short-range 1LL Yukawa potentials shown in Figure 4. Across all simulated LDOS results for both odd and even $N_{e}$ systems under various e-e interactions, the most experimentally detectable lowest peaks under the previous assumption correspond to states having very large relative angular momentum within the MR anyon cluster. However, depending on the strength of the bias voltage and the details of the STM tip, it does not necessarily correspond to the lowest energy state.
Notably, even for MR phases where the energy spectra appear similar under various realistic electron-electron interactions, as we mentioned in the last section, the LDOS remains sensitive to these variations, manifested from the strengths and locations of the density peaks in Figure 6. This underscores STM’s potential as a tool for detecting and verifying effective anyon interactions in both Abelian and non-Abelian FQH systems. However, compared to the Laughlin state—where peaks are well separated even at lower resolutions—MR systems require measurements of higher resolution (for example, FWHM = $0.001\ e^{2}/\epsilon l_{B}$ in Figure 6). This is understandable since removing one electron creates three Laughlin quasiholes, but four MR quasiholes, and the latter leads to many more configurations with similar energies.
We emphasize that while detailed peak positions and splittings in the LDOS depend on the microscopic form of the interaction, certain qualitative features are universal. For example, in the Laughlin case, the emergence of two dominant peak clusters from the 3-quasihole spectrum is present across different interaction models and FWHM values, where the lower cluster contributes to the quasihole excitations. In the Moore–Read case, the LDOS shows a clear distinction between even and odd electron number sectors, a direct consequence of the underlying non-Abelian fusion rules. These features are expected to survive under reasonable experimental conditions and provide a possible path toward distinguishing topological signatures from non-universal effects in STM experiments.
V Conclusion
In this paper, we investigated the effective interactions of anyons in Laughlin and Moore-Read FQH phases derived from diverse electron-electron interactions, both model and realistic ones. The effective interaction within anyons is an interpolation between fermions’ repulsive and bosons’ attractive force, exhibiting various and unusual behaviors depending on the detailed forms of the electron-electron interaction in the 2DEG. The recently developed STM experiments in FQH systems could be a new platform to explore these properties of anyons. As we show, many bound anyon states with different internal structures will be involved in the low-lying energy parts of the tunneling spectrum. We point out that effective anyon interactions are important factors in understanding this part of the results in STM tunneling experiments. Since anyons can be considered the elementary particles for the corresponding FQH phases and are potentially useful for the storage and processing of quantum information, the interplay between their effective interactions and the correlated topological phases should be considered as a significant feature when investigating their properties both theoretically and for future applications.
Note.
While this manuscript was under review, a related preprint appeared  [77] on the formation of anyon clusters (anyon “molecule”), readers may find it complementary.
Acknowledgements.
Some of the numerics in this work are performed using the DiagHam package for which we would like to thank all the authors. This work is supported by the National Research Foundation, Singapore, under the NRF Fellowship Award (NRF-NRFF12-2020-005), Singapore Ministry of Education (MOE) Academic Research Fund Tier 3 Grant (No. MOE-MOET32023-0003) “Quantum Geometric Advantage”, and Singapore Ministry of Education (MOE) Academic Research Fund Tier 2 Grant (No. MOE-T2EP50124-0017).